# From Majority to Minority: A Diffusion-based Augmentation for Underrepresented Groups in Skin Lesion Analysis

### v1

@@ 4 Results
This trend of consistent improvement is also evident in the more challenging scenarios where there are little or no reference images from the minority groups (Tables 3 and 4) in the training set. Notably, significant improvement is observed when just a few reference images from the minority group are available in the training set for image generation and classification. The transformer-based classifier demonstrates a larger improvement gap over the real image baseline than the CNN-based models. In the most challenging setting, with no reference images from the minority group, the improvement margin narrowed, suggesting our pipeline effectively maximizes the use of limited information from the seed images of the minority groups in the training process. Despite these challenges, the sustained improvements in the third setting validate our framework’s effectiveness in transferring information across groups. Examples of real and synthetic image pairs for each condition are presented in Fig. 2. Qualitatively, the synthetic images generated by our augmentation framework introduce more diversity to the training sets, including variations in skin color and lesion patterns.
@@
Architecture Train Accuracy Precision Recall F1 Score
VGG-16 real 55.58 $\pm$ 0.10 54.60 $\pm$ 0.11 51.84 $\pm$ 0.29 51.97 $\pm$ 0.45
syn 57.62 $\pm$ 0.09 56.62 $\pm$ 0.15 55.42 $\pm$ 0.20 55.36 $\pm$ 0.22
real + syn 58.08 $\pm$ 0.08 57.36 $\pm$ 0.13 55.78 $\pm$ 0.17 55.85 $\pm$ 0.23
ResNet-18 real 49.42 $\pm$ 0.36 49.79 $\pm$ 0.44 48.30 $\pm$ 0.23 47.42 $\pm$ 0.30
syn 53.47 $\pm$ 0.32 52.39 $\pm$ 0.30 53.79 $\pm$ 0.30 51.97 $\pm$ 0.29
real + syn 56.50 $\pm$ 0.15 55.72 $\pm$ 0.15 55.10 $\pm$ 0.16 54.76 $\pm$ 0.18
ViT-B-16 real 57.66 $\pm$ 0.35 61.89 $\pm$ 0.68 55.45 $\pm$ 0.36 55.96 $\pm$ 0.51
syn 58.94 $\pm$ 0.23 62.79 $\pm$ 0.20 55.62 $\pm$ 0.22 56.53 $\pm$ 0.18
real + syn 60.96 $\pm$ 0.31 64.57 $\pm$ 0.01 57.51 $\pm$ 0.29 58.92 $\pm$ 0.31
@@
Table 4: Classification results when no image from the minority group is in the training set. The test set has 291 non-seed images of the minority group (FST V-VI).
@@
To further investigate our framework’s generation capabilities, we conduct an ablation study comparing our framework with various generation strategies. This study focuses on the first setting, where the test set consists of light- or dark-skinned seed images (56 for each type). We first examine the Stable Diffusion’s vanilla text-to-image and image-to-image pipelines to generate synthetic images. Next, we leverage Textual Inversion to learn the lesion embeddings and generate synthetic images from these embeddings, with text-to-image and image-to-image pipelines. Since image-to-image outperforms text-to-image in both vanilla SD and Textual Inversion generation, we focus on image-to-image generation after fine-tuning the diffusion model using LoRA, to investigate if optimizing the diffusion model can benefit the generation even more. We train a VGG-16 using only the synthetic images and then compare these generation strategies with ours, as shown in Table 5.
@@
Overall, training classifiers on synthetic images generated by text-to-image models proves less effective than employing image-to-image techniques, underscoring the importance of visual cues in augmenting skin lesion classification. Additionally, using off-the-shelf models for image generation yields less improvement than utilizing training strategies such as Textual Inversion and LoRA, regardless of whether the target is a minority or majority group. Finally, the combination of Textual Inversion and LoRA results in the highest accuracy, thereby validating the practicality of our design which integrates these two strategies. This improvement can be explained by the model’s enhanced ability to associate the fine visual cues of the lesion with the learned textual tokens.
@@
Since a direct comparison with existing related works is challenging due to the uncertain use of data, this ablation study can serve as an indirect comparison. As introduced previously, related works either leveraged off-the-shelf diffusion models such as DALL·E or fine-tuned a Stable Diffusion model for text-to-image generation. Our results demonstrate that utilizing the dual guidance of visual cues and text via fine-tuning diffusion models can maximize the potential of diffusion-based augmentation to enhance the diagnosis for minority groups.
@@
Figure 2: Examples of real and synthetic image pairs. Synthetic images were generated by a model trained on light-skinned images only, using prompts conditioned on dark skin types.
@@
Table 5: Classification accuracy with various generation strategies for the first setting. Here, “vanilla” stands for Stable Diffusion’s original text-to-image and image-to-image pipelines, “ti” for Textual Inversion, and “lora” for LoRA.
@@ 5 Conclusion
In this work, we present an effective diffusion-based augmentation framework that consistently improves classification results for the minority group, even when training the classifier exclusively with synthetic images. This improvement is observed regardless of the availability of reference data from the minority group in the training set. The ablation study also validates that our framework’s dual-guidance generation approach successfully learns novel lesion concepts previously unfamiliar to the diffusion models. A practical takeaway from this study is that, even in cases of extreme data scarcity, existing data and diffusion models can still provide valuable insights, maximizing information usage and achieving better performance. In the future, we plan to apply this technique to other medical datasets characterized by significant differences in group sizes. Additionally, as we used all synthetic images generated for each setting without any filtering mechanism, we also aim to investigate which types of synthetic data are useful for lesion diagnosis and how to generate them.

### v2

@@ 4 Results
This trend of consistent improvement is also evident in the more challenging scenarios where there are little or no reference images from the minority groups (Tables 3 and 4) in the training set. Notably, significant improvement is observed when just a few reference images from the minority group are available in the training set for image generation and classification. The transformer-based classifier demonstrates a larger improvement gap over the real image baseline than the CNN-based models. In the most challenging setting, with no reference images from the minority group, the improvement margin narrowed, suggesting that our pipeline effectively maximizes the use of limited information from the flexible subset of the minority group during training. Despite these challenges, the sustained improvements in the third setting validate our framework’s effectiveness in transferring information across groups. Examples of real and synthetic image pairs for each condition are presented in Fig. 2. Qualitatively, the synthetic images generated by our augmentation framework introduce more diversity to the training sets, including variations in skin color and lesion patterns.
@@
     Architecture Train Type Train Size Accuracy Precision Recall F1 Score
VGG-16 real 1228 55.58 $\pm$ 0.10 54.60 $\pm$ 0.11 51.84 $\pm$ 0.29 51.97 $\pm$ 0.45
syn 1228 * 5 57.62 $\pm$ 0.09 56.62 $\pm$ 0.15 55.42 $\pm$ 0.20 55.36 $\pm$ 0.22
real + syn 1228 * 6 58.08 $\pm$ 0.08 57.36 $\pm$ 0.13 55.78 $\pm$ 0.17 55.85 $\pm$ 0.23
ResNet-18 real 1228 49.42 $\pm$ 0.36 49.79 $\pm$ 0.44 48.30 $\pm$ 0.23 47.42 $\pm$ 0.30
syn 1228 * 5 53.47 $\pm$ 0.32 52.39 $\pm$ 0.30 53.79 $\pm$ 0.30 51.97 $\pm$ 0.29
real + syn 1228 * 6 56.50 $\pm$ 0.15 55.72 $\pm$ 0.15 55.10 $\pm$ 0.16 54.76 $\pm$ 0.18
ViT-B-16 real 1228 57.66 $\pm$ 0.35 61.89 $\pm$ 0.68 55.45 $\pm$ 0.36 55.96 $\pm$ 0.51
syn 1228 * 5 58.94 $\pm$ 0.23 62.79 $\pm$ 0.20 55.62 $\pm$ 0.22 56.53 $\pm$ 0.18
real + syn 1228 * 6 60.96 $\pm$ 0.31 64.57 $\pm$ 0.01 57.51 $\pm$ 0.29 58.92 $\pm$ 0.31
@@
Table 4: Classification results when no image from the minority group is in the training set, which only has non-flexible images of the majority group (1228 of FST I-II). The test set has non-flexible images of the minority group (291 of FST V-VI).
@@
To further investigate our framework’s generation capabilities, we conduct an ablation study comparing our framework with various generation strategies. This study focuses on the first setting, where the test set consists of light- or dark-skinned flexible images (56 for each type). We first examine the Stable Diffusion’s vanilla text-to-image and image-to-image pipelines to generate synthetic images. Next, we leverage Textual Inversion to learn the lesion embeddings and then generate synthetic images from these embeddings, with text-to-image and image-to-image pipelines. Since image-to-image outperforms text-to-image in both vanilla SD and Textual Inversion generation, we focus on image-to-image generation after fine-tuning the diffusion model using LoRA, to investigate if optimizing the diffusion model can benefit the generation even more. We train a VGG-16 using only the synthetic images and then compare these generation strategies with ours, as shown in Table 5.
@@
Overall, training classifiers on synthetic images generated by text-to-image models proves less effective than employing image-to-image techniques, underscoring the importance of visual cues in augmenting skin lesion classification. Additionally, using off-the-shelf models for image generation yields less improvement than training strategies such as Textual Inversion and LoRA, regardless of whether the target is a minority or majority group. Finally, the combination of Textual Inversion and LoRA results in the highest accuracy, thereby validating the practicality of our design which integrates these two strategies. This improvement can be explained by the model’s enhanced ability to associate the fine visual cues of the lesion with the learned textual tokens.
@@
Since a direct comparison with existing related works is challenging due to the uncertain use of data, this ablation study can serve as an indirect comparison. As introduced previously, related works leveraged off-the-shelf diffusion models such as DALL·E or fine-tuned a Stable Diffusion model for text-to-image generation. Our results demonstrate that utilizing the dual guidance of visual cues and text prompts via fine-tuning diffusion models can maximize the potential of diffusion-based augmentation and enhance the diagnosis for minority groups.
@@
Figure 2: Examples of synthetic images generated by a model trained exclusively on light-skinned images, using prompts describing dark skin types.
@@
Table 5: Classification accuracy with various generation strategies for the first setting. Here, “vanilla" stands for Stable Diffusion’s original text-to-image and image-to-image pipelines, “ti" for Textual Inversion, and “lora" for LoRA.
@@ 5 Conclusion
In this work, we present an effective diffusion-based augmentation framework that consistently improves classification results for the minority group, even when training the classifier exclusively with synthetic images. This improvement is observed regardless of the availability of reference data from the minority group in the training set. The ablation study also validates that our framework’s dual-guidance generation approach successfully learns novel lesion concepts previously unfamiliar to the diffusion models. A practical takeaway from this study is that, even in cases of data scarcity, existing data and diffusion models can still provide valuable insights, maximizing information usage and achieving better performance. In the future, we plan to apply this technique to other medical datasets characterized by significant differences in group sizes. Additionally, as we used all synthetic images generated for each setting without any filtering mechanism, we also aim to investigate which types of synthetic data are useful for lesion diagnosis and how to generate them.
{credits}
5.0.1 Acknowledgements
This work was partly supported by the NSF EPSCoR-Louisiana Materials Design Alliance (LAMDA) program #OIA-1946231 and partly by the Harold L. and Heather E. Jurist Center of Excellence for Artificial Intelligence at Tulane University.